# Functors on the category of finite sets revisited

### v2

@@ 4.1. An explicit idempotent
For $n\in\mathbb{N}$ and a $\mathbbm{k}\mathbf{FA}$-module $F$, one has:
@@
$\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n},F)\cong\mathrm{image}\Big{(}F(\mathbf{n+1})\stackrel{{\scriptstyle F(\pi_{\mathbf{n}})}}{{\longrightarrow}}F(\mathbf{n+1})\Big{)}\subset F(\mathbf{n+1}).$
4.2. Morphisms between tensor products of $\overline{P}$
Corollary 4.3 allows us to calculate $\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(\overline{P}^{\otimes s},\overline{P}^{\otimes t})$ for $s\geq t$.
@@ 4.2. Morphisms between tensor products of $\overline{P}$
For $s,t\in\mathbb{N}$, if $s>t$ then $\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(\overline{P}^{\otimes s},\overline{P}^{\otimes t})=0$. For $s=t$, the place permutation action on $\overline{P}^{\otimes t}$ induces an isomorphism $\mathbbm{k}\mathfrak{S}_{t}\cong\mathrm{End}_{\mathcal{F}(\mathbf{FA})}(\overline{P}^{\otimes t})$ of $\mathbbm{k}$-algebras.
@@
The case $t=0$ is treated directly as follows. If $s>0$, then $\overline{P}^{\otimes s}(\mathbf{1})=0$, hence the image of any map $\ \overline{P}^{\otimes s}\rightarrow\overline{P}^{\otimes 0}=\overline{\mathbbm{k}}$ must be a proper sub-functor of $\overline{\mathbbm{k}}$. Since $\overline{\mathbbm{k}}$ is simple (as is easily checked), the image is zero. The case $s=0=t$ is immediate.
Now consider the case $t>0$. By the short exact sequence (3.4), the key step is to analyse $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n},\overline{P}^{\otimes t})$. We show that this vanishes if $n>t$ (so that $n>0$), by applying Corollary 4.3 with $F=\overline{P}^{\otimes t}$. (For notational simplicity in this proof, for a map $g$ of $\mathbbm{k}\mathbf{FA}$, we write $g$ rather than $F(g)$.)
For any finite set $X$ and chosen element $x\in X$, by Lemma 3.2 $\overline{P}^{\otimes t}(X)$ has a (non-canonical) basis given by the elements
(4.1) $\bigotimes_{i=1}^{t}([y_{i}]-[x])$
where $(y_{i})\in(X\backslash\{x\})^{\times t}$. In particular, this applies with $X=\mathbf{n+1}$, taking $x=1$.
Consider applying the map $[f_{Y}]$ to the basis element (4.1), with $Y\subset\{2,\ldots,n+1\}\subset\mathbf{n+1}$. Clearly, if some $y_{i}$ belongs to $Y$, then the basis element maps to zero; otherwise $[f_{Y}]$ acts by the identity on this basis element. Thus, to calculate the image of the basis element (4.1) under $\pi_{\mathbf{n}}$, we only need consider the $Y$ such that $Y\cap\bigcup_{i=1}^{t}\{y_{i}\}=\emptyset$. Explicitly, the idempotent $\pi_{\mathbf{n}}$ acts on this basis element by multiplication by
$\sum_{Y\subset\{2,\ldots n+1\}\backslash\bigcup_{i=1}^{t}\{y_{i}\}}(-1)^{|Y|}.$
A straightforward calculation shows that this sum is zero if $\{2,\ldots n+1\}\backslash\bigcup_{i=1}^{t}\{y_{i}\}$ is non-empty and is $1$ otherwise. This corresponds to $\pi_{\mathbf{n}}$ acting on this basis element by 00 in the first case and by the identity in the second.
Clearly, if $n>t$, then $\bigcup_{i=1}^{t}\{y_{i}\}\subsetneq\{2,\ldots,n+1\}$ for any $(y_{i})\in Y^{\times t}$. Thus, by the above, Corollary 4.3 gives
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n},\overline{P}^{\otimes t})=0\mbox{\quad if $n>t$.}$
Using the projection $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\twoheadrightarrow\overline{P}^{\otimes n}$ induced by $P^{\mathbf{FA}}_{\mathbf{1}}\twoheadrightarrow\overline{\mathbbm{k}}$, this implies that
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},\overline{P}^{\otimes t})=0\mbox{\quad if $n>t$.}$
This argument adapts to treat the case $n=t$. Indeed, using the short exact sequence (3.4) and the above, the projection $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}\twoheadrightarrow\overline{P}^{\otimes t}$ induces an isomorphism
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes t},\overline{P}^{\otimes t})\cong\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t},\overline{P}^{\otimes t})$
and the right hand side can be calculated as above. Namely, $\pi_{\mathbf{t}}$ acts non-trivially on the basis element (4.1) if and only if $i\mapsto y_{i}$ is a bijection from $\mathbf{t}$ to $\{2,\ldots,t+1\}$ and, in this case, $\pi_{\mathbf{t}}$ acts as the identity.
The map $\mathbbm{k}\mathfrak{S}_{t}\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes t},\overline{P}^{\otimes t})$ given by place permutations of the tensor factors is clearly an injective map of $\mathbbm{k}$-algebras. The above argument shows that it is an isomorphism of the underlying $\mathbbm{k}$-vector spaces, by comparing dimensions. ∎
@@
    The calculation in the case $s\leq t$ is much more delicate. This was first carried out in [PV23]; a direct approach is given in Part III working over a field of characteristic zero (see Theorem 11.9).
@@
    Theorem 4.4 is our substitute for the squishing arguments employed by Wiltshire-Gordon in [Wil14]. These approaches are clearly related.
4.3. Finite length
We record the following general consequence of Corollary 3.11:
Proposition 4.6.
Suppose that $F$ is a simple $\mathbbm{k}\mathbf{FA}$-module, then either $F=\mathbbm{k}_{\mathbf{0}}$ or there exists $n\in\mathbb{N}$ such that $F$ is a quotient of $\overline{P}^{\otimes n}$. Moreover, if $F$ is a composition factor of $\overline{P}^{\otimes t}$, then $F(\mathbf{t+1})\neq 0$.
@@ 4.3. Finite length
By Corollary 3.11, if $F$ is a simple $\mathbbm{k}\mathbf{FA}$-module, then $F$ is a quotient either of $\mathbbm{k}$ or of a functor of the form $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$, for some $n\in\mathbb{N}$. In the first case, one must have $F=\mathbbm{k}_{\mathbf{0}}$. In the second, using the short exact sequence (3.4), either $F$ is a quotient of $\overline{P}^{\otimes n+1}$ or of $\overline{P}^{\otimes n}$. This yields the first statement.
Now suppose that $F$ is a composition factor of $\overline{P}^{\otimes t}$. As above, there is a surjection $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\twoheadrightarrow F$ for some $n\in\mathbb{N}$ (the case $F=\mathbbm{k}_{\mathbf{0}}$ is excluded, since $\overline{P}^{\otimes t}(\mathbf{0})=0$). Since $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n+1}}$, by Yoneda, it follows that $F(\mathbf{n+1})\neq 0$.
Now, since $F$ is a composition factor of $\overline{P}^{\otimes t}$ by hypothesis and $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$ is projective, the surjection to $F$ lifts to a non-zero map $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\rightarrow\overline{P}^{\otimes t}$. Theorem 4.4 therefore implies that $n\leq t$. The fact that $F(\mathbf{n+1})$ is non-zero thus implies that $F(\mathbf{t+1})\neq 0$, as required, since $\mathrm{Id}_{\mathbf{n+1}}$ factors across $\mathbf{t+1}$ in $\mathbf{FA}$. ∎
From Proposition 4.6, one deduces:
@@
For $n\in\mathbb{N}$, the $\mathbbm{k}\mathbf{FA}$-module $\overline{P}^{\otimes n}$ has finite length (i.e., has a finite length composition series). Hence the indecomposable projectives of $\mathcal{F}(\mathbf{FA})$ have finite length.
@@
The functor $\overline{P}^{\otimes n}$ takes finite-dimensional values, since it is a subfunctor of $P^{\mathbf{FA}}_{\mathbf{n}}$, which takes finite-dimensional values. Now, by Proposition 4.6, each composition factor of $\overline{P}^{\otimes n}$ contributes non-trivially to $\overline{P}^{\otimes n}(\mathbf{n+1})$. Since the latter is a finite dimensional $\mathbbm{k}$-vector space, the first statement follows.
An indecomposable projective of $\mathcal{F}(\mathbf{FA})$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$ for some $n\in\mathbb{N}$. Using the short exact sequence (3.4), one has that $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$ has finite length. ∎
4.4. A model for the right augmentation to $\mathbbm{k}{\mathbf{FS}}$
Recall from Proposition 2.11 that, for $n\in\mathbb{N}$, there is a surjection
(4.2) $\mathbbm{k}\mathbf{FA}(-,\mathbf{n})\twoheadrightarrow\mathbbm{k}{\mathbf{FS}}(-,\mathbf{n})$
of $\mathbbm{k}\mathfrak{S}_{n}\otimes\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules. The purpose of this section is to explain how this is related to the canonical inclusion
(4.3) $\overline{P}^{\otimes n}\hookrightarrow P^{\mathbf{FA}}_{\mathbf{n}},$
considered as a morphism of $\mathbbm{k}\mathfrak{S}_{n}^{\mathrm{op}}\otimes\mathbbm{k}\mathbf{FA}$-modules. (Recall that $\overline{P}^{\otimes 0}$ is $\overline{\mathbbm{k}}$, by convention.)
Theorem 4.8.

### v3

@@ 4.1. Splitting $P^{\mathbf{FA}}_{X}$ using Möbius inversion
For $(X,x)$ a finite pointed set, there is a direct sum decomposition:
$P^{\mathbf{FA}}_{X}\cong\bigoplus_{\emptyset\subseteq Y\subseteq X\backslash\{x\}}\mathbf{Q}^{X\backslash Y,x}_{\emptyset}.$
Hence the set of objects $\{P^{\mathbf{FA}}_{\mathbf{0}}=\mathbbm{k}\}\amalg\{\mathbf{Q}^{\mathbf{n},1}_{\emptyset}\mid 0<n\in\mathbb{N}\}$ is a set of projective generators of $\mathcal{F}(\mathbf{FA})$.
Proof.
The first statement follows from (4.1) by using Proposition 4.7. The second statement follows by using the fact that $\{P^{\mathbf{FA}}_{\mathbf{n}}\mid n\in\mathbb{N}\}$ is a set of projective generators of $\mathcal{F}(\mathbf{FA})$. ∎
Remark 4.9.
In terms of the identification provided by Proposition 4.14 below, this result was proved in an earlier version of this paper by a direct, ad hoc proof, using only the results of Section 3. The current proof makes clear how the result relates to Möbius inversion, thus making the result more conceptual.
Corollary 4.8 has a counterpart for the decomposition of the standard projectives in $\mathcal{F}(\mathbf{FA}^{\mathrm{op}})$ using the following projectives:
Definition 4.10.
For $(X,x)$ a finite pointed set and $T\subseteq X\backslash\{x\}$, define $\mathbf{R}^{X,x}_{T}$ by
$\mathbf{R}^{X,x}_{T}:=e^{X,x}_{T}\mathbbm{k}\mathbf{FA}(\bullet,X)\in\mathrm{Ob}\hskip 2.0pt\mathcal{F}(\mathbf{FA}^{\mathrm{op}}).$
In the following, $\mathbbm{k}_{\mathbf{0}}$ is the $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-module supported on $\emptyset$ with value $\mathbbm{k}$.
Corollary 4.11.
For $(X,x)$ a finite pointed set, there is a direct sum decomposition of projective objects in $\mathcal{F}(\mathbf{FA}^{\mathrm{op}})$:
$P^{\mathbf{FA}^{\mathrm{op}}}_{X}\cong\bigoplus_{\emptyset\subseteq Y\subseteq X\backslash\{x\}}\mathbf{R}^{X\backslash Y,x}_{\emptyset}.$
Hence, the set of objects $\{P^{\mathbf{FA}^{\mathrm{op}}}_{\emptyset}=\mathbbm{k}_{\mathbf{0}}\}\amalg\{\mathbf{R}^{\mathbf{n},1}_{\emptyset}\mid 0<n\in\mathbb{N}\}$ is a set of projective generators of $\mathcal{F}(\mathbf{FA}^{\mathrm{op}})$.
Remark 4.12.
For $(X,x)$ a finite pointed set and $T\subseteq X\backslash\{x\}$, the functors $\mathbf{Q}^{X,x}_{T}$ and $\mathbf{R}^{X,x}_{T}$ are related by using Yoneda’s lemma, as follows. Consider $\mathbbm{k}\mathbf{FA}=\mathbbm{k}\mathbf{FA}(-,-)$ as a bifunctor, i.e., an object of $\mathcal{F}(\mathbf{FA}^{\mathrm{op}}\times\mathbf{FA})$, then there are isomorphisms:
$\mathbf{R}^{X,x}_{T}$ $\cong$ $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}^{X,x}_{T},\mathbbm{k}\mathbf{FA})$
$\mathbf{Q}^{X,x}_{T}$ $\cong$ $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}^{\mathrm{op}}}(\mathbf{R}^{X,x}_{T},\mathbbm{k}\mathbf{FA}).$
4.2. Identifying the functors $\mathbf{Q}^{(X,x)}_{\emptyset}$ and $\mathbf{R}^{(X,x)}_{\emptyset}$
Fix $(X,x)$ a finite pointed set, so that $X$ is non-empty. Using tensor products indexed by finite sets, there is a canonical isomorphism $P^{\mathbf{FA}}_{X}\cong(P^{\mathbf{FA}}_{\mathbf{1}})^{\otimes X}$; using the distinguished $x\in X$, the latter can be rewritten as $P^{\mathbf{FA}}_{\mathbf{1}}\otimes(P^{\mathbf{FA}}_{\mathbf{1}})^{\otimes X\backslash\{x\}}$.
Lemma 4.13.
For $(X,x)$ a finite pointed set, there is a surjection
@@ 4.2. Identifying the functors $\mathbf{Q}^{(X,x)}_{\emptyset}$ and $\mathbf{R}^{(X,x)}_{\emptyset}$
$P^{\mathbf{FA}}_{X}\cong P^{\mathbf{FA}}_{\mathbf{1}}\otimes(P^{\mathbf{FA}}_{\mathbf{1}})^{\otimes X\backslash\{x\}}$ $\stackrel{{\scriptstyle\rho^{X,x}}}{{\twoheadrightarrow}}$ $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes X\backslash\{x\}}$
$\ [y]\otimes\bigotimes_{j\in X\backslash\{x\}}[y_{j}]$ $\mapsto$ $[y]\otimes\bigotimes_{j\in X\backslash\{x\}}([y_{j}]-[y])$
where $y_{j}]-[y]$ is considered as a section of $\overline{P}$ via Lemma 3.3.
For a finite set $Y$ and $\varphi\in\mathbf{FA}(X,Y)$, $\varphi]$ lies in the kernel of $\rho^{X,x}(Y)$ if and only if $\varphi^{-1}\varphi(x)\supsetneq\{x\}$.
Proof.
The morphism $\rho^{X,x}$ can be constructed by using Yoneda’s lemma. The fact that it is surjective follows from Lemma 3.3.
The second statement follows from a direct calculation. ∎
Proposition 4.14.
For $(X,x)$ a finite pointed set, the surjection $\rho^{X,x}$ induces an isomorphism
$\mathbf{Q}^{X,x}_{\emptyset}\cong P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes X\backslash\{x\}}.$
[…]
Proof.
The first statement follows from the observation that the composite
$P^{\mathbf{FA}}_{X}\cong P^{\mathbf{FA}}_{\mathbf{1}}\otimes(P^{\mathbf{FA}}_{\mathbf{1}})^{\otimes X\backslash\{x\}}\stackrel{{\scriptstyle\rho^{X,x}}}{{\twoheadrightarrow}}P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes X\backslash\{x\}}\hookrightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes(P^{\mathbf{FA}}_{\mathbf{1}})^{\otimes X\backslash\{x\}}\cong P^{\mathbf{FA}}_{X}$
(considered as an element of $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{X},P^{\mathbf{FA}}_{X})\cong\mathbbm{k}\mathbf{FA}(X,X)$) identifies with $e^{X,x}_{\emptyset}$, by inspection using the first part of Lemma 4.13. The second statement then follows from the second statement of that Lemma. ∎
Remark 4.15.
The second part of Proposition 4.14 allows one to give an explicit description of the action of morphisms of $\mathbf{FA}$ on $\mathbf{Q}^{X,x}_{\emptyset}$. (Similarly, a description of the structure of the $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-module $\mathbf{R}^{X,x}_{\emptyset}$ can be derived from Proposition 4.17 below.)
Corollary 4.16.
For $(X,x)$ a finite pointed set, there is a short exact sequence
(4.2) $0\rightarrow\overline{P}^{\otimes X}\rightarrow\mathbf{Q}^{X,x}_{\emptyset}\rightarrow\overline{P}^{\otimes X\backslash\{x\}}\rightarrow 0$
[…]
Proof.
By construction, one has the short exact sequence
$0\rightarrow\overline{P}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\overline{\mathbbm{k}}\rightarrow 0.$
Upon forming the tensor product with $\overline{P}^{\otimes X\backslash\{x\}}$ one obtains the required short exact sequence, by using the isomorphism of Proposition 4.14. ∎
One can also analyse the projective $\mathbf{R}^{X,x}_{\emptyset}$, considered as a subfunctor of $P^{\mathbf{FA}^{\mathrm{op}}}_{X}\cong\mathbbm{k}\mathbf{FA}(\bullet,X)$.
Proposition 4.17.
For $(X,x)$ a finite pointed set and $Y$ a finite set, $\mathbf{R}^{X,x}_{\emptyset}(Y)$ is a free, finite rank $\mathbbm{k}$-submodule of $\mathbbm{k}\mathbf{FA}(Y,X)$ with basis $e^{X,x}_{\emptyset}[\psi]$, where $\psi\in\mathbf{FA}(Y,X)$ ranges over the set of maps such that $\psi(Y)\cup\{x\}=X$.
Proof.
One first checks that, if $\psi(Y)\cup\{x\}\subsetneq X$, then $e^{X,x}_{\emptyset}[\psi]=0$. Indeed, by definition, one has
$e^{X,x}_{\emptyset}[\psi]=\sum_{T\subseteq X\backslash\{x\}}(-1)^{|T|}[f^{X,x}_{T}\circ\psi].$
[…]
It remains to check that the elements $e^{X,x}_{\emptyset}[\psi]$, where $\psi\in\mathbf{FA}(Y,X)$ ranges over the set of maps such that $\psi(Y)\cup\{x\}=X$, are linearly independent. For such a $\psi$, consider $\psi_{T}:=f^{X,x}_{T}\circ\psi\in\mathbf{FA}(Y,X)$, for $T\subseteq X\backslash\{x\}$. By definition of $f^{X,x}_{T}$, if $T\neq\emptyset$, $\exists t\in T$ such that $t\not\in\mathrm{image}f^{X,x}_{T}$, with $t\neq x$, so that $\psi_{T}(Y)\cup\{x\}\subsetneq X$. Writing
$e^{X,x}_{\emptyset}[\psi]=[\psi]+\sum_{T\neq\emptyset}(-1)^{|T|}[\psi_{T}],$
and using that $\mathbf{FA}(Y,X)$ is a basis for $\mathbbm{k}\mathbf{FA}(Y,X)$, the result follows easily by neglecting all terms $\varphi]$ such that $\varphi(Y)\cup\{x\}\subsetneq X$. ∎
Corollary 4.18.
For $(X,x)$ a finite pointed set, there is a short exact sequence in $\mathcal{F}(\mathbf{FA}^{\mathrm{op}})$:
[…]
In particular, for a finite set $Y$, $\mathbf{R}^{X,x}_{\emptyset}(Y)=0$ if $|X|\geq|Y|+2$.
Proof.
[…]
It remains to check that this corresponds to a short exact sequence of functors, as stated. This follows by checking that the map $\mathbbm{k}{\mathbf{FS}}(Y,X\backslash\{x\})\rightarrow R^{X,x}_{\emptyset}=e^{X,x}_{\emptyset}\mathbbm{k}\mathbf{FA}(Y,X)$ given by $\psi]\mapsto e^{X,x}_{\emptyset}[\psi]$ defines a natural inclusion. This is straightforward. ∎
4.3. Morphisms between functors of the form $\mathbf{Q}_{\emptyset}^{X,x}$
The main result of this section, Theorem 4.19, gives information on morphisms between projectives of the form $\mathbf{Q}_{\emptyset}^{X,x}$. It was proposed by the referee as a more general and more conceptual result from which the author’s original Corollary 4.21 can be derived.
@@ 4.3. Morphisms between functors of the form $\mathbf{Q}_{\emptyset}^{X,x}$
For finite pointed sets $(X,x)$ and $(Y,y)$, there are natural isomorphisms:
[…]
@@
By Yoneda, we have the identification
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}^{X,x}_{\emptyset},\mathbf{Q}^{Y,y}_{\emptyset})=e^{X,x}_{\emptyset}\mathbbm{k}\mathbf{FA}(Y,X)e^{Y,y}_{\emptyset}$
and the right hand side can be rewritten as $\mathbf{R}^{X,x}_{\emptyset}(Y)e^{Y,y}_{\emptyset}$, a submodule of $\mathbf{R}^{X,x}_{\emptyset}(Y)$.
The vanishing for $|X|\geq|Y|+2$ follows from that given in Corollary 4.18.
[…]
As above, $\mathbf{R}^{X,x}_{\emptyset}(Y)$ identifies with $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}^{X,x}_{\emptyset},P^{\mathbf{FA}}_{Y})$. Now, by Corollary 4.8, there is a direct sum decomposition
$P^{\mathbf{FA}}_{Y}\cong\bigoplus_{\emptyset\subseteq Z\subseteq Y\backslash\{y\}}\mathbf{Q}^{Y\backslash Z,y}_{\emptyset}.$
[…]
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}^{X,x}_{\emptyset},\mathbf{Q}^{Y,y}_{\emptyset})\cong\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}^{X,x}_{\emptyset},P^{\mathbf{FA}}_{Y})$
(this is independent of $y\in Y$). The result follows. ∎
Example 4.20.
[…]
Moreover, by inspection, the dotted map factors as
$\mathbf{Q}^{X,x}_{\emptyset}\twoheadrightarrow\overline{P}^{\otimes X\backslash\{x\}}\hookrightarrow\mathbf{Q}^{X\backslash\{x\},y}_{\emptyset}.$
This yields an explicit construction of the maps appearing in the short exact sequence (4.2) of Corollary 4.16.
Corollary 4.21.
@@
For $s,t\in\mathbb{N}$, if $s>t$ then $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes s},\overline{P}^{\otimes t})=0$. For $s=t$, the place permutation action on $\overline{P}^{\otimes t}$ induces an isomorphism of $\mathbbm{k}$-algebras: $\mathbbm{k}\mathfrak{S}_{t}\cong\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes t})$.
Proof.
Using the morphisms appearing in (4.2) (as reconstructed in Example 4.20), one has the morphisms $\mathbf{Q}^{\mathbf{s+1},s+1}_{\emptyset}\twoheadrightarrow\overline{P}^{\otimes s}$ and $\overline{P}^{\otimes t}\hookrightarrow\mathbf{Q}^{\mathbf{t},t}_{\emptyset}$. These induce the inclusion
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes s},\overline{P}^{\otimes t})\hookrightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}^{\mathbf{s+1},s+1}_{\emptyset},\mathbf{Q}^{\mathbf{t},t}_{\emptyset}).$
This immediately implies the vanishing for $s>t$, using the vanishing statement of Theorem 4.19.
The case $s=t$ follows by verifying that the composite $\mathbbm{k}\mathfrak{S}_{t}\hookrightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes t},\overline{P}^{\otimes t})\hookrightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}^{\mathbf{t+1},t+1}_{\emptyset},\mathbf{Q}^{\mathbf{t},t}_{\emptyset})$, where the first map is given by the place permutation action, is the isomorphism given by Theorem 4.19. This is straightforward. ∎
@@
    The calculation in the case $s<t$ is much more delicate. This was first carried out in [7]; a direct approach is given in Part III working over a field of characteristic zero (see Theorem 11.9).
@@
    Theorem 4.19 and Corollary 4.21 are our substitutes for the squishing arguments employed by Wiltshire-Gordon in [10].
4.4. Calculating $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$ and applications
For $n\in\mathbb{N}$, we consider $\overline{P}^{\otimes n}$ (using our convention that, for $n=0$, this is $\overline{\mathbbm{k}}$); the symmetric group $\mathfrak{S}_{n}$ acts on $\overline{P}^{\otimes n}$ by place permutations.
Putting together Corollary 4.16 with the identification given in Example 4.20, we have a projective presentation of $\overline{P}^{\otimes n}$:
(4.3) $\mathbf{Q}^{\mathbf{n+2},n+2}_{\emptyset}\rightarrow\mathbf{Q}^{\mathbf{n+1},n+1}_{\emptyset}\rightarrow\overline{P}^{\otimes n}\rightarrow 0.$
The first map is induced by the canonical inclusion $\mathbf{n+1}\subset\mathbf{n+2}$, as in Example 4.20.
Using this presentation, we prove the following:
Theorem 4.23.
For $n\in\mathbb{N}$, there is an isomorphism of $\mathbbm{k}\mathfrak{S}_{n}\otimes\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})\cong\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n}),$
where the $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-module structure of the codomain is given by Proposition 2.11 and its $\mathbbm{k}\mathfrak{S}_{n}$-module structure by the action of $\mathfrak{S}_{n}$ on $\mathbf{n}$.
@@
Applying the functor $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,P^{\mathbf{FA}}_{\bullet})$ to the presentation (4.3) gives the exact sequence
$0\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})\rightarrow\mathbf{R}^{\mathbf{n+1},n+1}_{\emptyset}\rightarrow\mathbf{R}^{\mathbf{n+2},n+2}_{\emptyset}.$
Here the structure of the two right hand terms is described by Proposition 4.17, with complements in Corollary 4.18. The morphism between these is induced by the inclusion $\mathbf{n+1}\subset\mathbf{n+2}$. It follows that this factors as
(4.4) $\mathbf{R}^{\mathbf{n+1},n+1}_{\emptyset}\twoheadrightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n+1})\hookrightarrow\mathbf{R}^{\mathbf{n+2},n+2}_{\emptyset},$
where the surjection and inclusion are those appearing in the short exact sequence of Corollary 4.18 for $(\mathbf{n+1},n+1)$ and $(\mathbf{n+2},n+2)$ respectively.
Hence, by the short exact sequence of that Corollary, one obtains the required isomorphism of $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules. One checks directly that this is $\mathbbm{k}\mathfrak{S}_{n}$-equivariant, as required. ∎
The argument shows more:
@@
For $n\in\mathbb{N}$, applying the functor $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,P^{\mathbf{FA}}_{\bullet})$ to the short exact sequence
$0\rightarrow\overline{P}^{\otimes n+1}\rightarrow\mathbf{Q}^{\mathbf{n+1},n+1}_{\emptyset}\rightarrow\overline{P}^{\otimes n}\rightarrow 0$
of Corollary 4.16 yields the short exact sequence
$0\rightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n})\rightarrow\mathbf{R}_{\emptyset}^{\mathbf{n+1},n+1}\rightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n+1})\rightarrow 0$
of Corollary 4.18.
@@
Using Theorem 4.23 and the left exactness of $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,P^{\mathbf{FA}}_{\bullet})$, the only point to establish is that the inclusion $\overline{P}^{\otimes n+1}\hookrightarrow\mathbf{Q}^{\mathbf{n+1},n+1}_{\emptyset}$ induces the surjection $\mathbf{R}_{\emptyset}^{\mathbf{n+1},n+1}\twoheadrightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n+1})$. This follows from the factorization given in (4.4). ∎
Corollary 4.25.